OpenStudy (kendricklamar2014):

Solve this system by graphing. Tell whether the system has one solution, infinitely many solutions or no solutions. {3x + y = 2 {4y = 12 - 12x

OpenStudy (kendricklamar2014):

@phi

OpenStudy (phi):

First step, in the second equation, rewrite it in standard form. that means add +12x to both sides what do you get ?

OpenStudy (kendricklamar2014):

Like this @Phi: 4y + 12x = 12 - 12x + 12x

TheSmartOne (thesmartone):

She meant standard form Ax+By=C So 4y = 12 - 12x +12x +12x -----------------

TheSmartOne (thesmartone):

And since it says graphing we can easily plug it into https://www.desmos.com/calculator And see.

OpenStudy (kendricklamar2014):

How do you add 12x to 4y??

TheSmartOne (thesmartone):

And when you see the graph. If the two lines are parallel then the lines have no solution If the two lines are the same then there are infinitely many solutions If the two lines intersect at one point then there is only one solution.

TheSmartOne (thesmartone):

Do you get that? @KendrickLamar2014

OpenStudy (kendricklamar2014):

So, the answer is no Solution @TheSmartOne

TheSmartOne (thesmartone):

You plug the 2 equations into the calculator.

OpenStudy (kendricklamar2014):

yes, and they were parallel

TheSmartOne (thesmartone):

Exactly!

OpenStudy (kendricklamar2014):

Thanks